A Seller sells his product on 14% profit of the market price. If the selling price is 3420 then find the
market price. 

Grade:12

2 Answers

Fauzan
4 years ago
Let the Market price be X.
A gain is seen in the given problem, hence,
Selling price = Market price + Profit
3420= X +14% of 3420
X= 3420 - 14% of 3420
X= 3420 - 478.8
X= 2941.2
 
 
 
Yuna Savanna
2 years ago
Capital = Php 2,000 expected increase after 3 years if its interest rate is 
 
a) 2% compounded monthly 
b) 3.5% compound quarterly
c) 7% compound annually